OpenStudy (satan312):

Which is an equation in point-slope form for the given point and slope? Point: ; Slope: A. B. C. D.

OpenStudy (satan312):

a y+4=-5/6(x+8) b y-4=-5/6(x+8) c y+4=-5/6(x-8) d y-4=-5/6(x-8)